For the years 2000–2005, the median price of a single-family home in the United States can be approximated by the exponential function A = 227,200(1.087)t, where t is the number of years after the year 2000. What is the growth rate of housing prices for that period? What is the price of the house in the year 2005?


&&&&&&&&&&&&&&&&&&&&&&&&&&&&&&&&&&&&&&&&&&&&&&&&&&&&&









Describe a situation that would be modeled by exponential growth. Write an exponential function to represent the growth.


Answer by stanbon(75887)

For the years 2000–2005, the median price of a single-family home in the United States can be approximated by the exponential function A = 227,200(1.087)t, where t is the number of years after the year 2000. What is the growth rate of housing prices for that period? What is the price of the house in the year 2005?
A(t) = 227,200(1.087)t
A(5) = 227,200(1.087)^5 = $344,791.10
---------------------------------------------------
Describe a situation that would be modeled by exponential growth. Write an exponential function to represent the growth.
-------
Compound interest investing is modeled by an exponential function.
$200 invested at 6% compounded quarterly is modeled as followed.
A(t) = 200(1 + (0.06/4))^(4t) where t is number of years in the future.
========================================================================
Cheers,
Stan H.